OpenStudy (anonymous):

(x-4/ x- 3) -( x-6/x^2 -3x) =1/5

OpenStudy (lalaly):

\[(\frac{x-4}{x-3} \times \frac{x}{x})-\frac{x-6}{x(x-3)}=\frac{1}{5}\]\[\frac{x(x-4)}{x(x-3)}-\frac{x-6}{x(x-3)}=\frac{1}{5}\]\[\frac{x^2-4x-x+6}{x(x-3)}=\frac{1}{5}\]\[\frac{x^2-5x+6}{x(x-3)}=\frac{1}{5}\]\[\frac{(x-3)(x-2)}{x(x-3)}=\frac{1}{5}\]\[\frac{x-2}{x}=\frac{1}{5}\]\[x=5(x-2)\]\[x=5x-10\]\[-4x=-10\]\[x=\frac{10}{4}\]